# Large Language Models for Recommendation with Deliberative User Preference Alignment

### v1

@@ Abstract.
While recent advancements in aligning Large Language Models (LLMs) with recommendation tasks have shown great potential and promising performance overall, these aligned recommendation LLMs still face challenges in complex scenarios. This is primarily due to the current alignment approach focusing on optimizing LLMs to generate user feedback directly, without incorporating deliberation. To overcome this limitation and develop more reliable LLMs for recommendations, we propose a new Deliberative Recommendation task, which incorporates explicit reasoning about user preferences as an additional alignment goal. We then introduce the Deliberative User Preference Alignment framework, designed to enhance reasoning capabilities by utilizing verbalized user feedback in a step-wise manner to tackle this task. The framework employs collaborative step-wise experts and tailored training strategies for each expert. Experimental results across three real-world datasets demonstrate the rationality of the deliberative task formulation and the superior performance of the proposed framework in improving both prediction accuracy and reasoning quality.
@@ 1. INTRODUCTION
In this light, we propose a Deliberative User Preference Alignment (DeliRec) framework to teach LLMs with the recognized reasoning capabilities with the guidance of verbalized user feedback (e.g., user reviews (Hasan et al., 2024)). As illustrated in Figure 2, DeliRec employs three collaborative experts, each dedicated to a specific reasoning step, to hone step-specific reasoning capabilities. To reduce the computational and memory cost, these experts are implemented using separate QLoRA (Dettmers et al., 2023) adapters on the same LLM. Recognizing that each reasoning step necessitates extracting distinct signals from the verbalized user feedback, each expert is equipped with a uniquely tailored training strategy. Taking rating prediction (Liu et al., 2019a) as an example, we conduct extensive experiments across three real-world datasets, validating the effectiveness of DeliRec regarding prediction accuracy and reasoning quality, and the rationality of the framework design. The code and datasets are available at https://anonymous.4open.science/r/OurDeliRec-F4B5.
@@
We propose a Deliberative User Preference Alignment framework, which achieves the reasoning process with step-wise experts associated with specifically designed training strategies.
We conduct extensive experiments on three datasets, validating the effectiveness and rationality of the proposed DeliRec framework, showing the potential of slow thinking in recommendation.
@@ 3. METHOD
In this section, we present DeliRec with three experts to collaboratively perform multi-step reasoning before recommendation. To supervise the reasoning process, DeliRec incorporates user reviews, directly aligning the reasoning with user preferences rather than only inferring preferences through users’ numerical or binary feedback (e.g., ratings and clicks). We then detail the three experts, i.e., summarizer, reasoner, and predictor, with their fine-tuning strategies. An overview of DeliRec is illustrated in Figure 2.
@@ 3.1. Summarizer for Preference Distillation
DeliRec leverages a summarizer to distill the user’s preferences and the item’s positive and negative features from their historical reviews. It filters out noise from the reviews, captures nuanced user preferences, and enriches item features with the LLM’s inherent knowledge. The distillation output is formatted as an aspect-preference summary, i.e., concise and representative keywords that encapsulate essential user preferences and item features, providing a reasoning foundation for the subsequent reasoning process.
@@ 3.2. Reasoner for Preference Matching
Given all the aspect-preference summaries of a user and an item, DeliRec adopts a reasoner to evaluate their matching degree, generating reasons for why the user might like or dislike the item.
@@ 3.3. Predictor for Feedback Prediction
DeliRec finally introduces a predictor for user feedback prediction. In this work, we investigate the classic rating prediction task (Zheng et al., 2017). The input for the predictor is the reasoning process generated by the summarizer and reasoner. In addition, considering that the user’s historical ratings serve as a valuable reference for rating prediction (Koren et al., 2009), we also include ratings as input. The user’s actual rating for the target item is used for supervision.
@@ 4. EXPERIMENT
RQ1: How does DeliRec perform compared to other baseline methods in terms of accuracy in predicting user feedback (i.e., ratings) and the quality of generated reasons?
RQ2: Why is the multi-step reasoning strategy in DeliRec essential, and how does it compare in effectiveness to one-step reasoning strategies?
@@ 4.1. Experimental Settings
To comprehensively evaluate performance, we compare DeliRec against three categories of recommendation methods: traditional methods (MF), review-based recommendation methods (DeepCoNN, NARRE, DAML), and LLM-based methods (GPT-4o, Rec-SAVER, EXP3TR).
@@
For review-based methods, we adopt the implementation555https://github.com/ShomyLiu/Neu-Review-Rec. from previous work (Liu et al., 2019b). For all reasoning-enhanced LLMRec methods (i.e., Rec-SAVER, EXP3RT and our DeliRec), we use GPT-3.5-turbo666https://platform.openai.com/docs/models#gpt-3-5-turbo. as the teacher model and fine-tune the LLama3-8B model777https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/. using the QLoRA technique. To accelerate the training and inference process, we leverage the Unsloth888https://github.com/unslothai/unsloth. framework.
@@ 4.2. Main Results (RQ1)
We compare our DeliRec method with baseline methods on the accuracy of user feedback (rating) prediction (Predictor output) and the quality of generated reasons (Reasoner output).
Accuracy of User Feedback Prediction. The comparison of user feedback prediction performance is summarized in Table 2. As shown in the table, for the MAE metric, our method outperforms all datasets, achieving an average MAE reduction of approximately 0.02 compared to the best baseline results. For the RMSE metric, our method achieves the best performance on the Amazon Music and Amazon Book datasets, while delivering comparable results to the best baseline on the Yelp dataset. These superior performances of DeliRec confirm that incorporating reasoning about user preference and item features before making the final prediction, i.e., applying deliberation, is effective in enhancing RecLLM performance, thereby validating the rationality of our formulation of deliberative recommendation. Notably, although Rec-SAVER and EXP3RT also incorporate reasoning, their reasoning is performed in a single step, unlike our approach, which involves multiple steps—preference distillation, matching, and prediction. This lack of true “deliberation” in a single step may lead to their relatively poorer performance.
Quality of Generated Reasons. We next analyze the quality of the reasons generated by the Reasoner in DeliRec, compared to two reasoning-aware baselines (i.e., Rec-SAVER and EXP3RT). Table 3 summarizes the results. As shown by the results of the two metrics, whether compared with target item reviews using the BLEURT metric or evaluated using the more powerful GPT model (corresponding to GPTScore metric), the reasons generated by our method significantly outperform those of the two reasoning-aware baselines. Our method achieves an average relative improvement of 7.9% on the GPTScore metric and 7.0% on the BLEURT metric across all datasets. The superiority of our method’s reasoning can be attributed to the fact that we teach LLMs with recognized reasoning capabilities, guided by verbalized user feedback (e.g., user reviews (Hasan et al., 2024)), while the baselines fail to leverage this guidance.
@@ 4.3. In-depth Analysis
In the following sections, we present an analysis of the impact of multi-step reasoning and verbalized preference feedback in DeliRec, followed by the evaluation of DeliRec’s inference cost. Lastly, we provide detailed cases comparing the reasoning generated by DeliRec with those from baseline methods to demonstrate its superior alignment with the user’s true preferences.
@@ 4.3.1. Impact of Multi-step Reasoning (RQ2)
To evaluate the impact of multi-step reasoning in DeliRec, we first conduct an ablation study to analyze the necessity of the component in each step (i.e., Summarizer, Reasoner and Predictor) and then assess the performance changes by replacing the multi-step reasoning mechanism with two alternative single-step reasoning strategies.
Necessity of Each Step. An ablation study on DeliRec is conducted by removing each step to assess its contribution. Specifically, we make the following changes in each experiment: 1) w/o Step 1 (i.e., Summarizer): Directly using user historical reviews to replace the aspect-preference summaries as input for Step 2 and Step 3; 2) w/o Step 2 (i.e., Reasoner): Remove the Reason Placeholder section from the Step 3 Prompt and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings; 3) w/o Step 3 (i.e., Predictor): Combine Step 2 and Step 3 into a single step, asking the LLMs to simultaneously generate the match analysis and predicted ratings based on the aspect-preference summaries and the user’s historical ratings.
@@
One-step v.s. Multi-step. To verify the effectiveness of the multi-step reasoning strategy in DeliRec, we compare it with two alternative strategies: (1) One-step reasoning, which generates both an entangled reasoning process and the ratings in a single step, and (2) Chain-of-Thought (CoT) reasoning, which generates the outputs of our three reasoning steps in one single prompt. We prepare the datasets separately based on the two above strategies and train the LLMs accordingly. In Figure 4, we summarize the results and make the following observations: 1) Both one-step reasoning and CoT reasoning strategies lead to a notable increase in prediction errors on both Music and Book datasets, as reflected by the higher MAE and RMSE values. 2) In contrast, the performance of the CoT reasoning strategy is inferior to that of the simpler one-step reasoning strategy. This may be due to the complexity of multiple tasks involved in the CoT reasoning strategy, which poses significant challenges for LLMs to effectively process simultaneously.
4.3.2. Effect of Reason Generation Approaches (RQ3)
To validate the effectiveness of our reason generation approach in Step 2 of DeliRec, we compare it against four alternative approaches: 1) Without reason: Remove Step 2 from DeliRec entirely and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings, as done in the ablation study; 2) Review-guided reasons: Use the target item reviews as direct supervision for the LLMs’ learning process in Step 2; 3) Review-inferred reasons: Use ChatGPT to generate reasons inferred from the target item reviews and employ these generated reasons as supervision for Step 2; 4) History-inferred reason: Use ChatGPT to generate reasons inferred from the aspect-preference summaries using the Step 2 Prompt and directly use them as supervision for Step 2 without applying filtering through our reward model.
@@ 4.3.2. Effect of Reason Generation Approaches (RQ3)
| DeliRec | 0.5442 | 0.7722 | 0.6029 | 0.8345 |
@@ 4.3.3. Effectiveness of Reward Model (RQ3)
4.3.4. Cost Analysis
To evaluate the efficiency of DeliRec, we compared its inference time and generation cost with those of reasoning-aware baselines (Rec-SAVER and EXP3RT). As shown in Table 6, despite dividing the reasoning process into three steps, DeliRec achieves comparable average inference time and generates a similar number of tokens as the baselines. Specifically, it requires approximately 6 seconds and generates an average of 148 tokens per prediction. This efficiency is primarily attributed to two factors: 1) the preference distillation step in DeliRec is preprocessed in advance, as discussed in Section 3.1, leaving only the last two steps to be executed during inference; and 2) DeliRec only decodes a single token during the final step, as explained in Section 3.3. These designs effectively minimize computation costs.
@@ 4.3.5. Case study
We conduct a case study to examine whether the reasons generated by DeliRec better align with true user preferences, as reflected in the target item reviews. We compare the reasoning results of DeliRec with those of EXP3RT and Rec-SAVER. Figure 5 illustrates an example. In this case, the target item review indicates that the user values the book for its practicality and helpfulness. DeliRec accurately identifies these features and aligns them with the user’s preferences, as highlighted in green. In contrast, EXP3RT incorrectly attributes the user’s preference to the storyline and characters. Rec-SAVER also fails to recognize both the item’s key features and the user’s actual preferences, despite predicting that the user would like the book.
@@ 5.1. LLM-based Recommendation
Some studies have explored fine-tuning LLMs to improve their reasoning ability specifically for RecLLMs (Kim et al., 2024; Tsai et al., 2024; Bismay et al., 2024; Wang et al., 2024d). For example, RecSAVER utilizes reasoning generated by a larger LLM as ground truth to fine-tune a smaller LLM, enabling the latter to learn reasoning patterns tailored for recommendation tasks (Tsai et al., 2024). EXP3RT (Kim et al., 2024) employs a similar fine-tuning strategy but introduces an additional preparatory step: to construct user and item profiles based on their review histories first. Leveraging these constructed profiles, EXP3RT reasons about user preferences and predicts possible ratings in the same single step. However, these fine-tuning methods face two main limitations. First, they compress the complex reasoning process into a single step and train it jointly, which makes it challenging for LLMs to effectively learn. In contrast, DeliRec decomposes the reasoning process into multiple steps, training each step independently for better performance. Second, they lack fine-grained supervision in training data selection, leading models to learn the reasoning patterns misaligned with users’ real preferences. DeliRec addresses this limitation by using verbalized user feedback as supervision, resulting in reasons that more accurately reflect users’ preferences and enhance the multi-step reasoning abilities of RecLLMs.
@@ 5.3. Review-based Recommendation
Review-based recommendation systems have been widely studied recently to address tasks such as rating prediction (Qiu et al., 2021; Zhao et al., 2019; Kim et al., 2016; Yang et al., 2023; Shuai et al., 2022). These methods typically adopt a dual tower architecture that employs a user encoder and an item encoder to capture the semantics of reviews as user and item embeddings. Then, the user and item embeddings are combined through a fusion layer to derive joint representations, which are passed to a rating prediction layer to produce the final output (Hasan et al., 2024). To obtain better embeddings of users and items from reviews, many advanced techniques have been explored. For instance, DeepCoNN (Zheng et al., 2017) employs convolutional neural networks (CNNs) to extract semantic information from historical reviews, achieving excellent results. Furthermore, attention mechanisms (Bahdanau et al., 2015) are also introduced at both the word and review levels to identify critical words and reviews that are most significant (Liu et al., 2019b; Tay et al., 2018; Wu et al., 2019). For example, NARRE (Chen et al., 2018) employs attention mechanisms to prioritize reviews that are most useful for improving embedding learning. While these methods leverage reviews to represent user preferences implicitly with embeddings, resulting in the lack of interoperability. In contrast, DeliRec generates the reasoning process for each step in explicit textual form, offering clearer and more interpretable explanations of user preferences. Additionally, most review-based recommendation methods primarily leverage reviews as input features for generating embeddings. However, DeliRec also utilizes reviews as supervisory signals on the output side, aligning the LLM’s reasoning with the user’s true preferences.

### v2

# Reason4Rec: Large Language Models for Recommendation with Deliberative User Preference Alignment

@@ Abstract.
While recent advancements in aligning Large Language Models (LLMs) with recommendation tasks have shown great potential and promising performance overall, these aligned recommendation LLMs still face challenges in complex scenarios. This is primarily due to the current alignment approach focusing on optimizing LLMs to generate user feedback directly, without incorporating deliberation. To overcome this limitation and develop more reliable LLMs for recommendations, we propose a new Deliberative Recommendation task, which incorporates explicit reasoning about user preferences as an additional alignment goal. We then introduce the Reasoning-powered Recommender framework for deliberative user preference alignment, designed to enhance reasoning capabilities by utilizing verbalized user feedback in a step-wise manner to tackle this task. The framework employs collaborative step-wise experts and tailored training strategies for each expert. Experimental results across three real-world datasets demonstrate the rationality of the deliberative task formulation and the superior performance of the proposed framework in improving both prediction accuracy and reasoning quality.
@@ 1. INTRODUCTION
In this light, we propose a Reasoning-powered Recommender (Reason4Rec) framework to teach LLMs with the recognized reasoning capabilities with the guidance of verbalized user feedback (e.g., user reviews (Hasan et al., 2024)). As illustrated in Figure 2, Reason4Rec employs three collaborative experts, each dedicated to a specific reasoning step, to hone step-specific reasoning capabilities. To reduce the computational and memory cost, these experts are implemented using separate QLoRA (Dettmers et al., 2023) adapters on the same LLM. Recognizing that each reasoning step necessitates extracting distinct signals from the verbalized user feedback, each expert is equipped with a uniquely tailored training strategy. Taking rating prediction (Liu et al., 2019a) as an example, we conduct extensive experiments across three real-world datasets, validating the effectiveness of Reason4Rec regarding prediction accuracy and reasoning quality, and the rationality of the framework design. The code and datasets are available at https://github.com/Peter-Fy/Reason4Rec.
@@
We propose a Reasoning-powered Recommender framework for deliberative user preference alignment, which achieves the reasoning process with step-wise experts associated with specifically designed training strategies.
We conduct extensive experiments on three datasets, validating the effectiveness and rationality of the proposed Reason4Rec framework, showing the potential of slow thinking in recommendation.
@@ 3. METHOD
In this section, we present Reason4Rec with three experts to collaboratively perform multi-step reasoning before recommendation. To supervise the reasoning process, Reason4Rec incorporates user reviews, directly aligning the reasoning with user preferences rather than only inferring preferences through users’ numerical or binary feedback (e.g., ratings and clicks). We then detail the three experts, i.e., summarizer, reasoner, and predictor, with their fine-tuning strategies. An overview of Reason4Rec is illustrated in Figure 2.
@@ 3.1. Summarizer for Preference Distillation
Reason4Rec leverages a summarizer to distill the user’s preferences and the item’s positive and negative features from their historical reviews. It filters out noise from the reviews, captures nuanced user preferences, and enriches item features with the LLM’s inherent knowledge. The distillation output is formatted as an aspect-preference summary, i.e., concise and representative keywords that encapsulate essential user preferences and item features, providing a reasoning foundation for the subsequent reasoning process.
@@ 3.2. Reasoner for Preference Matching
Given all the aspect-preference summaries of a user and an item, Reason4Rec adopts a reasoner to evaluate their matching degree, generating reasons for why the user might like or dislike the item.
@@ 3.3. Predictor for Feedback Prediction
Reason4Rec finally introduces a predictor for user feedback prediction. In this work, we investigate the classic rating prediction task (Zheng et al., 2017). The input for the predictor is the reasoning process generated by the summarizer and reasoner. In addition, considering that the user’s historical ratings serve as a valuable reference for rating prediction (Koren et al., 2009), we also include ratings as input. The user’s actual rating for the target item is used for supervision.
@@ 4. EXPERIMENT
RQ1: How does Reason4Rec perform compared to other baseline methods in terms of accuracy in predicting user feedback (i.e., ratings) and the quality of generated reasons?
RQ2: Why is the multi-step reasoning strategy in Reason4Rec essential, and how does it compare in effectiveness to one-step reasoning strategies?
@@ 4.1. Experimental Settings
To comprehensively evaluate performance, we compare Reason4Rec against three categories of recommendation methods: traditional methods (MF), review-based recommendation methods (DeepCoNN, NARRE, DAML), and LLM-based methods (GPT-4o, Rec-SAVER, EXP3TR).
@@
For review-based methods, we adopt the implementation555https://github.com/ShomyLiu/Neu-Review-Rec. from previous work (Liu et al., 2019b). For all reasoning-enhanced LLMRec methods (i.e., Rec-SAVER, EXP3RT and our Reason4Rec), we use GPT-3.5-turbo666https://platform.openai.com/docs/models#gpt-3-5-turbo. as the teacher model and fine-tune the LLama3-8B model777https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/. using the QLoRA technique. To accelerate the training and inference process, we leverage the Unsloth888https://github.com/unslothai/unsloth. framework.
@@ 4.2. Main Results (RQ1)
We compare our Reason4Rec method with baseline methods on the accuracy of user feedback (rating) prediction (Predictor output) and the quality of generated reasons (Reasoner output).
Accuracy of User Feedback Prediction. The comparison of user feedback prediction performance is summarized in Table 2. As shown in the table, for the MAE metric, our method outperforms all datasets, achieving an average MAE reduction of approximately 0.02 compared to the best baseline results. For the RMSE metric, our method achieves the best performance on the Amazon Music and Amazon Book datasets, while delivering comparable results to the best baseline on the Yelp dataset. These superior performances of Reason4Rec confirm that incorporating reasoning about user preference and item features before making the final prediction, i.e., applying deliberation, is effective in enhancing RecLLM performance, thereby validating the rationality of our formulation of deliberative recommendation. Notably, although Rec-SAVER and EXP3RT also incorporate reasoning, their reasoning is performed in a single step, unlike our approach, which involves multiple steps—preference distillation, matching, and prediction. This lack of true “deliberation” in a single step may lead to their relatively poorer performance.
Quality of Generated Reasons. We next analyze the quality of the reasons generated by the Reasoner in Reason4Rec, compared to two reasoning-aware baselines (i.e., Rec-SAVER and EXP3RT). Table 3 summarizes the results. As shown by the results of the two metrics, whether compared with target item reviews using the BLEURT metric or evaluated using the more powerful GPT model (corresponding to GPTScore metric), the reasons generated by our method significantly outperform those of the two reasoning-aware baselines. Our method achieves an average relative improvement of 7.9% on the GPTScore metric and 7.0% on the BLEURT metric across all datasets. The superiority of our method’s reasoning can be attributed to the fact that we teach LLMs with recognized reasoning capabilities, guided by verbalized user feedback (e.g., user reviews (Hasan et al., 2024)), while the baselines fail to leverage this guidance.
@@ 4.3. In-depth Analysis
In the following sections, we present an analysis of the impact of multi-step reasoning and verbalized preference feedback in Reason4Rec, followed by the evaluation of Reason4Rec’s inference cost. Lastly, we provide detailed cases comparing the reasoning generated by Reason4Rec with those from baseline methods to demonstrate its superior alignment with the user’s true preferences.
@@ 4.3.1. Impact of Multi-step Reasoning (RQ2)
To evaluate the impact of multi-step reasoning in Reason4Rec, we first conduct an ablation study to analyze the necessity of the component in each step (i.e., Summarizer, Reasoner and Predictor) and then assess the performance changes by replacing the multi-step reasoning mechanism with two alternative single-step reasoning strategies.
Necessity of Each Step. An ablation study on Reason4Rec is conducted by removing each step to assess its contribution. Specifically, we make the following changes in each experiment: 1) w/o Step 1 (i.e., Summarizer): Directly using user historical reviews to replace the aspect-preference summaries as input for Step 2 and Step 3; 2) w/o Step 2 (i.e., Reasoner): Remove the Reason Placeholder section from the Step 3 Prompt and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings; 3) w/o Step 3 (i.e., Predictor): Combine Step 2 and Step 3 into a single step, asking the LLMs to simultaneously generate the match analysis and predicted ratings based on the aspect-preference summaries and the user’s historical ratings.
@@
One-step v.s. Multi-step. To verify the effectiveness of the multi-step reasoning strategy in Reason4Rec, we compare it with two alternative strategies: (1) One-step reasoning, which generates both an entangled reasoning process and the ratings in a single step, and (2) Chain-of-Thought (CoT) reasoning, which generates the outputs of our three reasoning steps in one single prompt. We prepare the datasets separately based on the two above strategies and train the LLMs accordingly. In Figure 4, we summarize the results and make the following observations: 1) Both one-step reasoning and CoT reasoning strategies lead to a notable increase in prediction errors on both Music and Book datasets, as reflected by the higher MAE and RMSE values. 2) In contrast, the performance of the CoT reasoning strategy is inferior to that of the simpler one-step reasoning strategy. This may be due to the complexity of multiple tasks involved in the CoT reasoning strategy, which poses significant challenges for LLMs to effectively process simultaneously.
@@
| Reason4Rec | 0.5442 | 0.7722 | 0.6029 | 0.8345 |
4.3.2. Effect of Reason Generation Approaches (RQ3)
To validate the effectiveness of our reason generation approach in Step 2 of Reason4Rec, we compare it against four alternative approaches: 1) Without reason: Remove Step 2 from Reason4Rec entirely and instruct the LLMs to directly predict the rating based solely on the aspect-preference summaries and historical ratings, as done in the ablation study; 2) Review-guided reasons: Use the target item reviews as direct supervision for the LLMs’ learning process in Step 2; 3) Review-inferred reasons: Use ChatGPT to generate reasons inferred from the target item reviews and employ these generated reasons as supervision for Step 2; 4) History-inferred reason: Use ChatGPT to generate reasons inferred from the aspect-preference summaries using the Step 2 Prompt and directly use them as supervision for Step 2 without applying filtering through our reward model.
@@ 4.3.3. Effectiveness of Reward Model (RQ3)
4.3.4. Cost Analysis
To evaluate the efficiency of Reason4Rec, we compared its inference time and generation cost with those of reasoning-aware baselines (Rec-SAVER and EXP3RT). As shown in Table 6, despite dividing the reasoning process into three steps, Reason4Rec achieves comparable average inference time and generates a similar number of tokens as the baselines. Specifically, it requires approximately 6 seconds and generates an average of 148 tokens per prediction. This efficiency is primarily attributed to two factors: 1) the preference distillation step in Reason4Rec is preprocessed in advance, as discussed in Section 3.1, leaving only the last two steps to be executed during inference; and 2) Reason4Rec only decodes a single token during the final step, as explained in Section 3.3. These designs effectively minimize computation costs.
@@ 4.3.5. Case study
We conduct a case study to examine whether the reasons generated by Reason4Rec better align with true user preferences, as reflected in the target item reviews. We compare the reasoning results of Reason4Rec with those of EXP3RT and Rec-SAVER. Figure 5 illustrates an example. In this case, the target item review indicates that the user values the book for its practicality and helpfulness. Reason4Rec accurately identifies these features and aligns them with the user’s preferences, as highlighted in green. In contrast, EXP3RT incorrectly attributes the user’s preference to the storyline and characters. Rec-SAVER also fails to recognize both the item’s key features and the user’s actual preferences, despite predicting that the user would like the book.
@@ 5.1. LLM-based Recommendation
Some studies have explored fine-tuning LLMs to improve their reasoning ability specifically for RecLLMs (Kim et al., 2024; Tsai et al., 2024; Bismay et al., 2024; Wang et al., 2024d). For example, RecSAVER utilizes reasoning generated by a larger LLM as ground truth to fine-tune a smaller LLM, enabling the latter to learn reasoning patterns tailored for recommendation tasks (Tsai et al., 2024). EXP3RT (Kim et al., 2024) employs a similar fine-tuning strategy but introduces an additional preparatory step: to construct user and item profiles based on their review histories first. Leveraging these constructed profiles, EXP3RT reasons about user preferences and predicts possible ratings in the same single step. However, these fine-tuning methods face two main limitations. First, they compress the complex reasoning process into a single step and train it jointly, which makes it challenging for LLMs to effectively learn. In contrast, Reason4Rec decomposes the reasoning process into multiple steps, training each step independently for better performance. Second, they lack fine-grained supervision in training data selection, leading models to learn the reasoning patterns misaligned with users’ real preferences. Reason4Rec addresses this limitation by using verbalized user feedback as supervision, resulting in reasons that more accurately reflect users’ preferences and enhance the multi-step reasoning abilities of RecLLMs.
@@ 5.3. Review-based Recommendation
Review-based recommendation systems have been widely studied recently to address tasks such as rating prediction (Qiu et al., 2021; Zhao et al., 2019; Kim et al., 2016; Yang et al., 2023; Shuai et al., 2022). These methods typically adopt a dual tower architecture that employs a user encoder and an item encoder to capture the semantics of reviews as user and item embeddings. Then, the user and item embeddings are combined through a fusion layer to derive joint representations, which are passed to a rating prediction layer to produce the final output (Hasan et al., 2024). To obtain better embeddings of users and items from reviews, many advanced techniques have been explored. For instance, DeepCoNN (Zheng et al., 2017) employs convolutional neural networks (CNNs) to extract semantic information from historical reviews, achieving excellent results. Furthermore, attention mechanisms (Bahdanau et al., 2015) are also introduced at both the word and review levels to identify critical words and reviews that are most significant (Liu et al., 2019b; Tay et al., 2018; Wu et al., 2019). For example, NARRE (Chen et al., 2018) employs attention mechanisms to prioritize reviews that are most useful for improving embedding learning. While these methods leverage reviews to represent user preferences implicitly with embeddings, resulting in the lack of interoperability. In contrast, Reason4Rec generates the reasoning process for each step in explicit textual form, offering clearer and more interpretable explanations of user preferences. Additionally, most review-based recommendation methods primarily leverage reviews as input features for generating embeddings. However, Reason4Rec also utilizes reviews as supervisory signals on the output side, aligning the LLM’s reasoning with the user’s true preferences.